Solve by rearranging the equations first. Then use the multiplication principle and then use the elimination method:
3x=8y+11
x+6y-8=0

Answer by nerdybill(7384)

Solve by rearranging the equations first. Then use the multiplication principle and then use the elimination method:
3x=8y+11
x+6y-8=0
.
Rearranging equations:
3x-8y = 11
x+6y = 8
.
Multiply both sides of bottom equation by -3:
3x-8y = 11
-3x-18y = -24
.
Add both equations together:
3x-8y = 11
-3x-18y = -24
---------------
-26y = -13
y = -13/(-26)
y = 1/2
.
Substitute above into bottom equation:
x+6y = 8
x+6(1/2) = 8
x+3 = 8
x = 5
.
Solution: ordered pair (5,1/2)